# WikiGap: Promoting Epistemic Equity by Surfacing Knowledge Gaps Between English Wikipedia and other Language Editions

### v2

@@ 4.2. Tasks
We constructed one quiz per article topic using facts extracted by InfoGap and surfaced through the WikiGap interface. From the 30 multilingual facts identified for each article, we manually selected a subset of 10, ensuring a roughly equal number of facts from each of the three language editions—French (fr), Russian (ru), and Chinese (zh)—to maintain balanced representation.
We then prompted a large language model GPT-4o (OpenAI, 2024b) to generate a multiple-choice question with four answer choices from each fact. The questions were manually reviewed and edited to ensure clarity, factual correctness, and alignment with the presented content. In addition to selecting the correct answer, participants were asked to indicate the language edition in which they found the supporting fact.
Although all quiz facts were sourced from non-English Wikipedia editions, we note that some participants may have inferred the correct answer from contextual clues in the English article. Nonetheless, the quizzes were designed to evaluate whether participants engaged with multilingual content. If participants had guessed entirely at random, the expected accuracy would be 25%, given the four-choice format. Example questions are included in the Appendix A.
@@
We recorded each participant’s task completion time and quiz accuracy score under both conditions. Completion time was used to assess how efficiently users could locate and interpret relevant information, while accuracy scores reflected knowledge retention and comprehension. These performance measures allowed us to quantitatively evaluate the effectiveness of WikiGap compared to the standard Wikipedia interface.
@@ 4.3. Materials and Topic Assignment
Topic assignment was designed to minimize cultural confounds and familiarity‑driven performance differences. Specifically, the food topic chosen for each participant could not be strongly linked to the participant’s self‑reported cultural or ethnic background. To further reduce individual‑performance bias, each topic appears under every experimental condition with data from exactly four different participants.
@@ 4.4. Procedure
The lead investigator conducted the study remotely over a 1–1.5 hour recorded Zoom call. At the start, participants received a zip file containing the WikiGap Chrome extension and were instructed to install it on their Google Chrome browser. A short tutorial was provided by the researcher for both the treatment and control conditions. During the treatment condition, we began the round by introducing the WikiGap extension’s functionality and allowed participants to explore the extension on their own until they felt ready to proceed. In the control condition, we demonstrated how to access language versions on Wikipedia using ILL, and how to use Chrome’s Google Translate function to read the content in English. After each round, participants completed a System Usability Scale (SUS) after finishing the quiz. We used the SUS to measure ease-of-use and users’ preference on with or without the WikiGap extension. For the first twelve participants, the evaluation concluded with an open-ended questionnaire. The remaining nine participants participated in a semi-structured interview. The questionnaire and interviews both covered the following six areas: overall impressions of the system, presentation of multilingual information, perception of cultural impact, trust in the augmented facts, perceived reliability of Wikipedia after using WikiGap, and future inclination to explore multilingual content.
@@ 5. Findings
① Users Strongly Preferred WikiGap over the Default ILL (D1, D2, D4)
@@
Participants attributed their strong preference for WikiGap not only to its ability to surface knowledge disparities, but also to the various design elements (D1-D4) that supported the quick access and ease of reading. Many highlighted how having access to multilingual content directly within the English article – through translated facts in a sidebar (D2, D4) – lowered the barrier to engaging with other language editions. It helped users “absorb information without the language difference obstacle” (P8). Participants appreciated the sidebar’s structure and organization by language, which kept all cross-lingual facts “apparent within a single page” (P5), eliminating the need to navigate away and reducing friction in multilingual exploration. In contrast to ILL’s full-page switch model, participants noted that the sentence-level fact card interface (D4) made the information easier to process. As P11 remarked, “[WikiGap] presents sentences instead of long paragraphs from the Wikipedia page.”
Other visual elements also shaped participants’ perceptions of usability. Several users found the in-text highlights (D1) helpful in surfacing areas of missing information in the English article. However, reactions to the color scheme used for language cues were mixed. While some found the colors intuitive, others, like P19, felt that the red-green coding could unintentionally suggest correctness or error: “The red color is usually an error indicator for me. I understand you’re using red and green to organize different language results, but at first glance they can look like ‘right or wrong’.”
@@
WikiGap Enhanced Fact-Finding Performance and Cross-Cultural Learning via Passive Discovery and Active Exploration (D3, D5)
To evaluate the impact of WikiGap on users’ ability to retrieve and retain cross-lingual information, we measured task performance through a fact-finding quiz, using both accuracy and completion time as metrics. As shown in Figure 8, participants performed significantly better when using WikiGap, achieving higher quiz accuracy and completing the task more quickly than with the default Wikipedia interface. A paired-samples t-test showed that accuracy was significantly higher in the WikiGap condition ($M$ = 0.91, $SD$ = 0.09) than in the control condition ($M$ = 0.73, $SD$ = 0.19), t(20) = 4.75, p ¡ .001. Completion time also improved significantly, with users finishing the quiz faster using WikiGap ($M$ = 12.39 minutes, $SD$ = 5.78) compared to the control ($M$ = 20.84 minutes, $SD$ = 8.68), t(20) = -5.52, p ¡ .001. These results suggest that WikiGap not only facilitated access to relevant information but also improved users’ efficiency in locating and interpreting cross-lingual facts. WikiGap consistently improved user performance across all five culturally diverse topics. For a topic-by-topic breakdown of quiz accuracy and completion time, see Appendix Figure 9.
@@
These performance gains were reinforced by participants’ qualitative feedback, which highlighted the value of both passive discovery and active exploration features. The keyword-based search box (D5) enabled users to directly retrieve relevant facts, supporting goal-oriented behaviors and speeding up information retrieval. As P16 explained, “I don’t need to go to multiple pages. I can search in the WikiGap extension and find any information quickly. Using the WikiGap extension saved me a lot of time.” This targeted retrieval helped users answer questions more efficiently, contributing to the improved fact-finding performance observed in the quiz metrics. Complementing this, the language filter controls (D3) allowed users to personalize their exploration by prioritizing facts from languages culturally aligned with the topic. Several participants reported that they first filtered to the language they deemed most relevant, as they“will trust more when the language is directly aligned with the original country of the food.” (P1). Similarly, P7 remarked, “I would love to see how the Tagalog page talks about Philippine adobo. In fact, diverse cultural perspectives on food is more linked to the cultural ecosystem where the food is grounded.”. Together, D3 and D5 supported efficient, goal-driven access to cross-lingual content.
@@
WikiGap Raises Awareness of Multilingual Gaps and Motivates Cross-Lingual Exploration
@@
For this participant and others, WikiGap did not merely expose content gaps – it fundamentally altered their perception of Wikipedia as a trustworthy authority. The perceived comprehensiveness of LLMs, combined with their ability to synthesize content across sources, positioned them as a more efficient alternative – even if they lacked transparency. [^5] In practice, the assumption that multilingual LLMs such as ChatGPT and Gemini can synthesize information learned from web text in different languages is not entirely true; they often fail to retrieve facts learned in one language when prompted in another language (Goldman et al., 2025).
@@
Although WikiGap was primarily designed to support readers by surfacing factual cross-lingual content, a few participants – particularly those with prior editing experience – reflected on how it could assist with editorial tasks. These insights point to an emergent opportunity for WikiGap to support not only knowledge consumption but also knowledge curation on Wikipedia.
@@ 6. Discussion
In this work, we sought to combat the epistemic injustice (Ajmani et al., 2023a) posed by the dominance of the English Wikipedia project over other language editions. We did so by leveraging novel affordances brought by advances in fine-grained cross-linguistic comparison (Samir et al., 2024), enabling us to identify factual gaps in English Wikipedia and succinctly display them to readers. In doing so, we challenged the belief in the English-as-superset assumption, with minimal disruption to the traditional English Wikipedia reading experience. Beyond enabling readers to access a more complete set of factual information, our work thus educates readers about the composition of knowledge on multilingual Wikipedia through a novel form of sociotechnical interaction. Below, we discuss how English Wikipedia came to have a limited demographic representation in its editor base, yet an outsized global readership (Sec. 6.1); how the English-as-superset hypothesis still persists among readers, despite its debunking in prior seminal studies from over a decade ago (e.g., Hecht and Gergle, 2010) (Sec. 6.2); and why this epistemology is a byproduct of the subsuming culture of computationalism (Golumbia, 2009), which (in part) explains its stubborn persistence (Sec. 6.3).
@@ 6.1. Editor-Readership Asymmetry in English Wikipedia
While the status of English as a global lingua franca is undeniable (Schneider, 2019), this doesn’t imply that English speakers from around the world can contribute to English Wikipedia. The broader Wikipedia project is no doubt motivated by an ethos of participatory, democratic knowledge curation. But historical inequities kneecap this libertarian ideal (Shaw and Hargittai, 2018). Participation in the internet has been unequal, both within and across nation-states, with greater access among socially privileged individuals. Four years after English Wikipedia was launched, access to the internet for developing countries was still in the single digits, percentage-wise (Chen and Wellman, 2005). Within nation states, internet usage is clearly stratified across gender, socioeconomic class, race, and a stark urban-rural divide (Chen and Wellman, 2005), many of which have also been studied in the context of English Wikipedia (Johnson et al., 2016; Hargittai and Shaw, 2015; Adams et al., 2019). The inequities that were present during the establishment of English Wikipedia matter, as peer production processes are understood to ossify and concentrate power among a small group of early editors, resembling an oligarchy (Shaw and Hill, 2014). Although these disparities are well studied, Wikimedia has struggled to reduce these editor demographic skews. As of (2024), English Wikipedia contributions are still overwhelmingly contributed by male-identifying editors, and there is a remarkable lack of Black and Hispanic editors from within the United States (Wikimedia Foundation, 2024).
@@
It is important to note that English Wikipedia attracts a massive readership, even outside of the geographic Anglosophere. As Kumar (2017) notes (at the time), $20\%$ of English Wikipedia’s visitors were based in India, even more than readers from the United Kingdom. The composition of English Wikipedia’s editor base is thus hardly representative of the broader population of literate, internet-equipped English speakers. As acknowledged in responses to our open-ended surveys, readers believe in the knowledge of local authorities. Indeed, the capacity for volunteers to contribute and disseminate information from geographically-situated sources was seen as one of the primary benefits of Wikipedia (Sen et al., 2015). There is thus a strong impetus to achieve more equitable knowledge representation. WikiGap presents an important contribution towards achieving greater epistemic equity.
@@ 6.2. English-as-superset Hypothesis Lives On
Although there are well-documented structural barriers to contributing to English Wikipedia, its overwhelming dominance makes it appear as a superset of the information contained in all other language editions, rather than a necessarily biased and socio-geographically-positioned construction (Kumar, 2021). Hecht (2013) studied this perception of English Wikipedia, calling it the English-as-superset hypothesis, writing that the hypothesis is “commonly adopted in Wikipedia-based research projects and systems”. Practically, this has entailed designating English Wikipedia as the ground-truth language edition for assessing, for example, the correctness of question-answering systems (Asai et al., 2021). Although there is now greater appreciation among Wikipedia researchers of the cross-linguistic information diversity between Wikipedia language editions, there is little reason to expect the broader Wikipedia userbase outside of Wikipedia experts to be privy to these results.
As Hickman et al. (2021) documented in their cross-linguistic comparative analysis of articles pertaining to Kashmir, editors of Hindi and Urdu articles drew on English Wikipedia for their primary research to a large extent. Their editing process thus implies an underlying acceptance of the English-as-superset hypothesis, where English Wikipedia is perceived as neutral, objective, and, critically, complete. Our user studies also demonstrate the prevalence of the hypothesis, with some participants commenting on the hypothesis directly: “Before WikiGap, I always assumed that everything presented in English would also be visible in another language.” Thus, while Hecht and Gergle (2010) effectively critiqued the English-as-superset hypothesis $15$ years ago, its continued prevalence calls for greater institutional intervention to increase awareness among the broader Wikipedia userbase of the vast cross-linguistic diversity between language editions. Our research demonstrates that the WikiGap interface directly challenges the English-as-superset hypothesis, while also providing discernible benefits for information seeking. As such, adoption of our interface can challenge this problematic epistemic assumption.

### v3

@@ 4.2. Tasks
We constructed one quiz per article topic using facts extracted by InfoGap and surfaced through the WikiGap interface. From the 30 multilingual facts identified for each article, we manually selected a subset of 10, ensuring a roughly equal number of facts from each of the three language editions—French (fr), Russian (ru), and Chinese (zh)—to maintain balanced representation. For a fair comparison, we ensured that all quiz questions are answerable from InfoGap facts and by extension from the respective articles in French, Russian, and Chinese.
We then prompted a large language model GPT-4o (OpenAI, 2024b) to generate a multiple-choice question with four answer choices from each fact. The questions were manually reviewed and edited to ensure clarity, factual correctness, and alignment with the presented content. In addition to selecting the correct answer, participants were asked to indicate the language edition in which they found the supporting fact. Although all quiz facts were sourced from non-English Wikipedia editions, we note that some participants may have inferred the correct answer from contextual clues in the English article. Nonetheless, the quizzes were designed to evaluate whether participants engaged with multilingual content. If participants had guessed entirely at random, the expected accuracy would be 25%, given the four-choice format. Example questions are included in the Appendix A.
@@
We recorded each participant’s task completion time and quiz accuracy score under both conditions. Quiz accuracy captured how much multilingual knowledge participants were able to obtain, while completion time reflected how quickly they could locate and interpret the relevant information. Taken together, these two measures provide an indication of learning efficiency and allowed us to quantitatively compare the effectiveness of WikiGap with the standard Wikipedia interface.
@@ 4.3. Materials and Topic Assignment
Topic assignment was designed to minimize cultural confounds and familiarity‑driven performance differences. Specifically, the food topic chosen for each participant could not be strongly linked to the participant’s self‑reported cultural or ethnic background. To reduce individual-performance bias, each topic was assigned in every condition to exactly four different participants.
@@ 4.4. Procedure
The lead investigator conducted the study remotely over a 1–1.5 hour recorded Zoom call. At the start, participants received a zip file containing the WikiGap extension and were instructed to install it on their Google Chrome browser. A short tutorial was provided by the researcher for both the treatment and control conditions. During the treatment condition, we began the round by introducing the WikiGap’s functionality and allowed participants to explore the extension on their own until they felt ready to proceed. In the control condition, we demonstrated how to access language versions on Wikipedia using ILLs, and how to use Chrome’s Google Translate function to read the content in English. At the beginning of each round, one of the five topics was assigned to the participant according to the topic assignment rules. After completing each round, the participant completed a System Usability Scale (SUS) after finishing the quiz. We used the SUS to measure ease-of-use and users’ preference on with or without the WikiGap extension. To conclude, the first twelve participants completed an open-ended questionnaire, while the remaining nine took part in a semi-structured interview. Both instruments covered the same six areas: overall impressions of the system, presentation of multilingual information, perceived cultural impact, trust in the augmented facts, perceptions of Wikipedia’s reliability after using WikiGap, and future inclination to explore multilingual content (see Appendix B for the full list of questions).
@@ 5. Findings
①Users Strongly Preferred WikiGap over the Default ILL (D1, D2, D3)
@@
Participants attributed their strong preference for WikiGap not only to its ability to surface knowledge disparities, but also to the various design elements (D1-D3) that supported the quick access and ease of reading. Many highlighted how having access to multilingual content directly within the English article – through translated facts in a sidebar (D2, D3) – lowered the barrier to engaging with other language editions. It helped users “absorb information without the language difference obstacle” (P8). Participants appreciated the sidebar’s structure and organization by language, which kept all cross-lingual facts “apparent within a single page” (P5), eliminating the need to navigate away and reducing friction in multilingual exploration. In contrast to ILL’s full-page switch model, participants noted that the sentence-level fact card interface (D3) made the information easier to process. As P11 remarked, “[WikiGap] presents sentences instead of long paragraphs from the Wikipedia page.”
Other visual elements also shaped participants’ perceptions of usability. Several users found the in-text gap markers (D1) helpful in surfacing areas of missing information in the English article. However, reactions to the color scheme used for language cues were mixed. While some found the colors intuitive, others, like P19, felt that the red-green coding could unintentionally suggest “right or wrong”.
@@
WikiGap Enhanced Fact-Finding Performance and Cross-Cultural Learning via Passive Discovery and Active Exploration (D4, D5)
To evaluate the impact of WikiGap on users’ ability to retrieve and retain cross-lingual information, we measured task performance through a fact-finding quiz, using both accuracy and completion time as metrics. As shown in Figure 7, participants performed significantly better when using WikiGap, achieving higher quiz accuracy and completing the task more quickly than with the default Wikipedia interface. A paired-samples t-test showed that accuracy was significantly higher in the WikiGap condition ($M$ = 0.91, $SD$ = 0.09) than in the control condition ($M$ = 0.73, $SD$ = 0.19), t(20) = 4.75, p ¡ .001. Completion time also improved significantly, with users finishing the quiz faster using WikiGap ($M$ = 12.39 minutes, $SD$ = 5.78) compared to the control ($M$ = 20.84 minutes, $SD$ = 8.68), t(20) = -5.52, p ¡ .001.555Faster completion times in the WikiGap condition is partly expected, as the quiz questions were constructed from facts extracted by InfoGap. In real-world use, not all facts would be presented by WikiGap, and completion time gains may therefore be more modest. These results suggest that WikiGap not only facilitated access to relevant information but also improved users’ efficiency in locating and interpreting cross-lingual facts. WikiGap consistently improved user performance across all five culturally diverse topics. For a topic-by-topic breakdown of quiz accuracy and completion time, see Appendix Figure 8.
@@
These performance gains were reinforced by participants’ qualitative feedback, which highlighted the value of both passive discovery and active exploration features. The fact search (D5) enabled users to directly retrieve relevant facts, supporting goal-oriented behaviors and speeding up information retrieval. As P16 explained, “I don’t need to go to multiple pages. I can search in the WikiGap extension and find any information quickly. Using the WikiGap extension saved me a lot of time.” This targeted retrieval helped users answer questions more efficiently, contributing to the improved fact-finding performance observed in the quiz metrics. Complementing this, the language filter controls (D4) allowed users to personalize their exploration by prioritizing facts from languages culturally aligned with the topic. Several participants reported that they first filtered to the language they deemed most relevant, as they“will trust more when the language is directly aligned with the original country of the food.” (P1). Similarly, P7 remarked, “I would love to see how the Tagalog page talks about Philippine adobo. In fact, diverse cultural perspectives on food is more linked to the cultural ecosystem where the food is grounded.”. Together, D4 and D5 supported efficient, goal-driven access to cross-lingual content.
@@
WikiGap Raises Awareness of Multilingual Gaps and Motivates Exploration
@@
For this participant and others, WikiGap altered their perception of Wikipedia as a trustworthy authority. The perceived comprehensiveness of LLMs, combined with their ability to synthesize content across sources, positioned them as a more efficient alternative, even if they lacked transparency. [^6] In practice, the assumption that multilingual LLMs such as ChatGPT and Gemini can synthesize information learned from web text in different languages is not entirely true; they often fail to retrieve facts learned in one language when prompted in another language (Goldman et al., 2025).
@@
Although WikiGap was primarily designed to support readers by surfacing factual cross-lingual content, a few participants, particularly those with prior editing experience, reflected on how it could assist with editorial tasks. These insights point to an emergent opportunity for WikiGap to support not only knowledge consumption but also knowledge curation on Wikipedia.
@@ 6. Discussion
In this work, we sought to combat the epistemic injustice (Ajmani et al., 2023a) posed by the dominance of the English Wikipedia project over other language editions. We did so by leveraging novel affordances brought by advances in fine-grained cross-linguistic comparison (Samir et al., 2024), enabling us to identify factual gaps in English Wikipedia and succinctly display them to readers. By surfacing complementary facts through WikiGap, our system performs an infrastructural inversion (Star, 1999): it makes visible the hidden cross-lingual knowledge gaps while productively challenging the long-standing English-as-superset assumption. Importantly, this intervention is introduced with minimal disruption to the familiar English Wikipedia reading experience. Beyond enabling access to a more complete set of factual information, our work educates readers about how knowledge on multilingual Wikipedia is composed and contested. Through this sociotechnical interaction, WikiGap demonstrates how systems can foster more equitable collaboration in distributed knowledge production across linguistic boundaries (Ackerman, 2000), reshaping how knowledge is consumed and shared across communities. Below, we discuss how English Wikipedia came to have a limited demographic representation in its editor base, yet an outsized global readership (Sec. 6.1); how the English-as-superset hypothesis still persists among readers, despite its debunking in prior seminal studies from over a decade ago (e.g., Hecht and Gergle, 2010) (Sec. 6.2); and why this epistemology is a byproduct of the subsuming culture of computationalism (Golumbia, 2009), which (in part) explains its stubborn persistence (Sec. 6.3).
@@ 6.1. Editor-Readership Asymmetry in English Wikipedia
While English has undeniably achieved the status of a global lingua franca (Schneider, 2019), this does not mean that English speakers worldwide have had equal opportunity to contribute to English Wikipedia. Wikipedia as a whole is animated by an ethos of participatory, democratic knowledge curation. Yet from its inception, historical inequities have undermined this libertarian ideal (Shaw and Hargittai, 2018). Participation in the internet has been unequal, both within and across nation-states, with greater access among socially privileged individuals. Four years after English Wikipedia was launched, access to the internet for developing countries was still in the single digits, percentage-wise (Chen and Wellman, 2005). Within nation states, internet usage is clearly stratified across gender, socioeconomic class, race, and a stark urban-rural divide (Chen and Wellman, 2005), many of which have also been studied in the context of English Wikipedia (Johnson et al., 2016; Hargittai and Shaw, 2015; Adams et al., 2019). The inequities that were present during the establishment of English Wikipedia matter, as peer production processes are understood to ossify and concentrate power among a small group of early editors, resembling an oligarchy (Shaw and Hill, 2014). This editor–reader asymmetry exemplifies a socio-technical gap (Ackerman, 2000), in which the ideals of open participation clash with the realities of unequal collaboration in a large-scale knowledge infrastructure.
Although these disparities are well studied, Wikimedia has struggled to reduce these editor demographic skews. As of (2024b), English Wikipedia contributions are still overwhelmingly contributed by male-identifying editors, and there is a remarkable lack of Black and Hispanic editors from within the United States (Wikimedia Foundation, 2024b).
@@
It is important to note that English Wikipedia attracts a massive readership, even outside of the geographic Anglosophere. As Kumar (2017) notes (at the time), $20\%$ of English Wikipedia’s visitors were based in India, even more than readers from the United Kingdom. The composition of English Wikipedia’s editor base is thus hardly representative of the broader population of literate, internet-equipped English speakers. As acknowledged in responses to our open-ended surveys, readers believe in the knowledge of local authorities. As one participant noted: “diverse cultural perspectives on food is more linked to the cultural ecosystem where the food is grounded” (P7). Indeed, the capacity for volunteers to contribute and disseminate information from geographically-situated sources was seen as one of the primary benefits of Wikipedia (Sen et al., 2015), Yet prior work highlights that readers are not simply passive consumers but active participants with distinct roles and needs, and that their preferences do not always align with where editorial effort is concentrated. This production–consumption gap poses a fundamental challenge for distributed collaborative work: designing socio-technical processes of knowledge creation that account for readers as well as editors (Antin and Cheshire, 2010; Lehmann et al., 2014). There is thus a strong impetus to achieve more equitable knowledge representation. WikiGap presents an important contribution towards achieving greater epistemic equity.
@@ 6.2. English-as-superset Hypothesis Lives On
Although there are well-documented structural barriers to contributing to English Wikipedia, its overwhelming dominance makes it appear as a superset of the information contained in all other language editions, rather than a necessarily biased and socio-geographically-positioned construction (Kumar, 2021). Hecht (2013) studied this perception of English Wikipedia, calling it the English-as-superset hypothesis, writing that the hypothesis is “commonly adopted in Wikipedia-based research projects and systems”. Practically, this has entailed designating English Wikipedia as the ground-truth language edition for assessing, for example, the correctness of question-answering systems (Asai et al., 2021).
Although there is now greater appreciation among Wikipedia researchers of the cross-linguistic information diversity between Wikipedia language editions, there is little reason to expect the broader Wikipedia userbase outside of Wikipedia experts to be privy to these results. As Hickman et al. (2021) documented in their cross-linguistic comparative analysis of articles pertaining to Kashmir, editors of Hindi and Urdu articles drew on English Wikipedia for their primary research to a large extent. Their editing process thus implies an underlying acceptance of the English-as-superset hypothesis, where English Wikipedia is perceived as neutral, objective, and, critically, complete.
Our user studies also demonstrate the prevalence of the hypothesis, with some participants commenting on the hypothesis directly: “Before WikiGap, I always assumed that everything presented in English would also be visible in another language.” Thus, while Hecht and Gergle (2010) effectively critiqued the English-as-superset hypothesis $15$ years ago, its continued prevalence calls for greater institutional intervention to increase awareness among the broader Wikipedia userbase of the vast cross-linguistic diversity between language editions.
Our research demonstrates that the WikiGap interface directly challenges the English-as-superset perception as well as the invisibility problem: The sentence-level complementary facts presented by WikiGap function as boundary objects (Star and Griesemer, 1989), serving users across different linguistic communities by making the otherwise hidden contributions of diverse editor communities visible and facilitating understanding of the distributed nature of collaborative knowledge production. As such, adoption of our interface can challenge the problematic status quo.